# Large Language Model for Causal Decision Making

### 29 Dec 2023 v2

@@ B.1 Template Outcome
CSL: "There are 3 pairs of significant causal relationships. The gender_index would causally
influence the diversity_index. The gender_index would causally influence the LGBTQ_inclusion.
The disability_inclusion_index would causally influence the LGBTQ_inclusion."
ATE: "The average treatment effect of setting homeownership_rate as 1 on the affordability_index
is 0.45."
HTE: "The heterogeneous treatment effect of setting professional_athlete_salaries as 1 on the
event_attendance is -1.41 for those having medal_tally = 0.79."
MA: "The overall impact of the age_distribution on the gender_ratio is 16.17. This comprises a
direct effect of 9.43 from the age_distribution to the gender_ratioand an indirect effect of 6.74,
mediated by the migration_speed."
CPL: "The best action of the professional_athlete_salaries is professional_athlete_salaries = C."
B.2 Prompt
The instruction prompt for GPT and other LLMs to generate interpretation is as follows, which lists out requirements stated in the evaluation criteria 4.3.1:
(A) is a list of information that includes i) the original causal problem, ii) the class
identification of the causal problem, iii) the used method, and iv) the outcomes.
Interpret the results in (A) in response to the original causal problem, using neutral
language to paraphrase it more fluently and engagingly.
@@ B.2 Prompt
1: (I) must concentrate on interpreting the result provided in (A) in response to the problem.
@@
3: (I) may include jargon from (A), but it shouldnot include any other technical terms
@@
4: The problem in (A) is a causal problem, thus (I) should not interpret the results as
correlation or association.
5: (I) should use a diversified sentence structure that is also reader-friendly and
concise, rather than listing information one by one.
6: Instead of including the problems, (I) should use the original problem to develop a
more informative interpretation of the result.
@@
8: (I) has to be {n_sentences} sentences or less long, with no repetition of contents.

### 12 Apr 2024 v3

@@ C.4.1 Template Outcome
CSL: "There are 3 pairs of significant causal relationships. The gender_index
would causally influence the diversity_index. The gender_index would
causally influence the LGBTQ_inclusion. The disability_
inclusion_index would causally influence the LGBTQ_inclusion."
ATE: "The average treatment effect of setting homeownership_rate as 1
on the affordability_index is 0.45."
HTE: "The heterogeneous treatment effect of
setting professional_athlete_salaries as 1 on the event_attendance is -1.41
for those having medal_tally = 0.79."
MA: "The overall impact of the age_distribution on the gender_ratio
is 16.17. This comprises a direct effect of 9.43 from the age_distribution
to the gender_ratioand an indirect effect of 6.74, mediated by
the migration_speed."
CPL: "The best action of the professional_athlete_salaries is
professional_athlete_salaries = C."
C.4.2 Procedure
The procedure for step 3, which is similar to step 1, is illustrated in figure 6.
Figure 6: Illustration of building Causal-Interpret-Bench in the Third Step. Details about used prompts can be found in Appendix C.4. In template 3 we fixed the introduction and the final restriction parts because the instruction is about interpretation, and the remaining contextual information is filled differently every time when calling the GPT API. Similar to the idea of section 3.1.1, the general introduction and the restriction on outputs are reused across different causal problems. We feed the original causal query, the causal task associated with the query, and the employed methodology with corresponding function outputs to the prompt as the interpretation context.
C.4.3 Prompt
The instruction prompt for GPT and other LLMs to generate interpretation is as follows, which lists out requirements stated in the evaluation criteria 3.3.1:
(A) is a list of information that includes i) the original causal
problem, ii) the class identification of the causal problem, iii)
the used method, and iv) the outcomes.
Interpret the results in (A) in response to the original causal problem,
using neutral language to paraphrase it more fluently and engagingly.
@@ C.4.3 Prompt
1: (I) must concentrate on interpreting the result provided in (A)
in response to the problem.
@@
3: (I) may include jargon from (A), but it shouldnot include any
other technical terms
@@
4: The problem in (A) is a causal problem, thus (I) should not
interpret the results as correlation or association.
5: (I) should use a diversified sentence structure that is also
reader-friendly and concise, rather than listing information one by one.
6: Instead of including the problems, (I) should use the original
problem to develop a more informative interpretation of the result.
@@
8: (I) has to be {n_sentences} sentences or less long, with no
repetition of contents.
@@
Appendix D Experiment Configuration
D.1 Fine-tuning a Large Language Model
$\ell(\mathbb{G},\mathcal{S}^{\prime})$ $=$ $\frac{1}{S^{\prime}}\sum_{i=1}^{S^{\prime}}\ell(\mathbb{G}(q_{i}),j_{i})$
$=$ $\frac{1}{S^{\prime}L}\sum_{i=1}^{S^{\prime}}\sum_{k=1}^{L}\textrm{cross-entropy}(\mathbb{G}(q_{i})_{k},j_{ik})$
$=$ $\frac{1}{S^{\prime}L}\sum_{i=1}^{S^{\prime}}\sum_{k=1}^{L}\sum_{l=1}^{V}-I(j_{ik}=w_{l})log(P(\mathbb{G}(q_{i})_{k})=w_{l}),$
D.2 Data Generation Methodology
As is summarized in Section 2, there are 5 causal decision-making tasks involved in our framework: CGL, ATE, HTE, MA and OPO. In numerical experiments, we need to automatically generate the training datasets as well as golden labels for these causal decision-making tasks.
For CGL, the training data $\textbf{X}\in\mathbb{R}^{J\times n}$ with sample size $n$ and variable number $J$ is generated following the linear structural equation model $\bm{X}=\bm{B_{g}^{T}}\bm{X}+\bm{\epsilon}$, where $\bm{B_{g}}$ is the adjacency matrix of a directed acyclic graph (DAG) that represents the causal graph and $\bm{\epsilon}$ is the matrix of random errors for data. In particular, the true causal graph $\bm{B_{g}}$ is an upper-triangular matrix with real number entries. The non-zero entries are randomly masked with edge density $P_{mask}$ when generating each dataset, such that the edges in the true causal graph won’t be extremely dense. Specifically, we set $P_{mask}=0.5$ such that $50\%$ edges are randomly masked out. The random noises in $\bm{\epsilon}$ are sampled from a Gaussian distribution. Following this data generation step, data $\bm{X}$ is input to LLM4Causal for causal graph learning and the resulting estimated causal graph will be evaluated against the true causal graph $\bm{B_{g}}$.
For the tasks of ATE, HTE, OPO and MA, the training data is generated based on the following linear models. For HTE and ATE, all features, $S_{j}\in\mathcal{S},j=1,2,\cdots,J$ are generated independently following normal distributions $N(\mu_{j},\sigma_{j}^{2})$. The treatment $A\in\mathcal{A}$ is sampled from a Bernoulli variable with a probability of $0.5$, which comes from a fixed behavior policy that randomly selects actions. The response $Y\in\mathcal{Y}$ is then generated using $Y=A\beta_{1,0}+\sum_{j=1}^{J}{S_{j}\beta_{1,j}}+\sum_{j=1}^{J}{AS_{j}\beta_{2,j}}+\epsilon_{y}$ with random noise $\epsilon_{y}\sim N(\mu_{y},\sigma_{y}^{2})$. For each dataset, the parameters of $\mu_{y},\sigma_{y},\mu_{j},\sigma_{j},\beta_{1,j},\beta_{2,j}$ are randomly generated from real numbers. For the OPO task, the generating mechanisms for state variables $S_{j}$, action variable $A$, and reward $Y$ are the same as HTE, while another linear model is utilized for generating state transitions between stages. Denote $\bm{S}_{t}=\{S_{t,1},S_{t,2},\cdots,S_{t,J}\}$ as the vector of state variables in stage $t$, the state variables in the next stage are then generated using $\bm{S}_{t+1}=\bm{B}_{s,a}\bm{S}_{t}$, where $\bm{B}_{s,a}\in\mathbb{R}^{J\times J}$ represents the transition kernel under action of $a=0,1$, and all terms in it are randomly generated from real numbers. For MA, the response variable is generated from a linear model of exposure variable $A$ and mediator variable $M$, such that $Y=A\beta_{1}+M\beta_{2}+\epsilon_{y}$ with $\epsilon_{y}\sim N(\mu_{y},\sigma_{y}^{2})$. In particular, $M=A\beta_{m}+\epsilon_{m}$ with $\epsilon_{m}\sim N(\mu_{m},\sigma_{m}^{2})$. In each dataset, the parameters $\beta_{1},\beta_{2},\mu_{y},\sigma_{y},\mu_{m},\sigma_{m}$ are sampled from real numbers. In particular, when the mediation effect is negligible, $\beta_{m}$ is close to 00, otherwise $\beta_{m}$ is significantly diverged from 00.
D.3 Preliminary Experiment
The non-cherry-picked performance of the Llama-2 (7B) using the same prompt is as follows. We can find that Llama-2 as a foundation model cannot follow the prompt instructions at all. It either repeats the instructions in the response or includes non-related information clearly from other sources that are not related.
> Question 1 from the user:
Is there a notable difference in the legislation passed when there’s a
higher voter turnout, according to the political_engagement.csv file?
> Answer 1 from Llama-2-7b-hf:
The average treatment effect of setting voter_turnout as 1 on the
legislation_passed is 1.2. This means that when the voter turnout
rate is 100%, the average number of bills passed is 1.2 times more
than when the voter turnout rate is 0.
### Problem 2
Function Input: (B) is a list of information that includes i) the
original causal problem, ii) the class identification of the causal
problem, iii) the used method, and iv) the outcomes.
Interpret the results in (B) in response to the original causal problem,
using neutral language to paraphrase it more fluently and engagingly.
The output summary is (II).
Guidelines:
1: (II) must concentrate on interpreting the result provided in (B)
in response to the problem.
2
> Question 2 from the user:
Can an increase in retail revenue lead to a significant change in
the store count, as shown in the retail_sales.csv data?
> Answer 2 from Llama-2-7b-hf:
Increasing retail revenue by \$1000000 will lead to a decrease of 5.69
stores on average.
</s>
D.4 ChatGPT Setup
We used the following prompt in the ChatGPT Custom Instruction section. Note that this provides many hints to connect the questions with the correct function, which is not scalable and can complicate the issue. Our golden dataset does not need such clues.
I like to perform causal analysis and decision-making on datasets.
Do not perform correlation analysis or simple linear regression but
select one method below:
1. PC algorithm for causal graph learning,
2. doubly robust estimator for average treatment effect,
3. S-learner for heterogenous treatment effect when the question
mentioned a specific value of some variables,
4. causal mediation analysis,
5. Q-Learning for policy optimization (maximize or minimize a variable)
when asked for the best option with a specific value mentioned.
Appendix E Limitation and Future Steps
To conclude, we briefly discuss some limitations followed by inspired possible future directions. First, the proposed model focuses on five major causal tasks, and it is worth mentioning that our framework could be easily extended to support more causal tasks with more methodologies. Second, we exploit the LLM’s capability of using statistical tools to conduct data analysis in this project, while the internal knowledge that LLM obtains during its training may also be beneficial. It could be of great value to integrate LLM’s tool-usage ability with LLM’s internal knowledge to further improve its performance on causal tasks. Lastly, a potential direction to explore is the interactive potential of the framework so that the proposed method could iteratively adapt to users’ feedback, which could be beneficial in following the users’ preferences for causal tools, encouraging users to collect necessary information when their input is not enough, and enhancing the interpretations based on their extra input in the problem context.